A sample of hydrogen gas has a pressure of 5.9 atm and a temperature of 290 K. After being cooled to temperature 200 K, what is the new pressure?


1
Expert's answer
2021-06-16T05:32:36-0400

P/T=k

5.9 / 290 = P2 / 200

P2 = 4.1 atm